OpenStudy (anonymous):

Help please! sin(-90˚)

OpenStudy (ghazi):

well, sin (-theta)= - sin theta, therefore sin (-90) = -0.8939 or approximately -1

OpenStudy (anonymous):

but how do you know it is 1 without a calculator

OpenStudy (ghazi):

well you have to calculate the value by this triangle, i assume that triangle of sides 1 unit and when you find any sine angle in that triangle and you use the limits in that you will get it around 0.89 or approximately 1, but when sine is -ve your value will be -1

OpenStudy (anonymous):

i dont understand this. how is there a triangle where the hypotenuse is 1?

OpenStudy (ghazi):

well if you just talk about a triangle where hypotenuse could be one , then you can use a triangle where sides are \[\frac{ 1 }{ \sqrt2 }\] but in this case consider an example  when theta 2= theta 1 and z=y , we have assumed y = x and when z=y then ratio of y/z will be =y/y and hence sin theta =1